OpenStudy (cicilybailey14):

Help with math stufff! Find the slope of the line passing through the given points. If the line is horizontal or vertical, state that. (-4, -5), (-3, -9)

OpenStudy (cicilybailey14):

@MakaylaTracy

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (-4,-5)...x1 = -4 and y1 = -5 (-3,-9)...x2 = -3 and y2 = -9 now we sub slope = (-9 - (-5) / (-3 - (-4) can you finish and find the slope...pay attention to the signs

OpenStudy (igreen):

Use the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) Plug in the two points: \(m = \dfrac{-9+5}{-3+4}\) Simplify it

OpenStudy (cicilybailey14):

-4

OpenStudy (igreen):

Yep, so the slope of the line is -4.

OpenStudy (texaschic101):

m = (-9 + 5) / (-3 + 4) = -4/1 = -4...yep...thats it one more thing that might help you... if you look at your 2 sets of points....if the x values were the same, then it is a vertical line. If the y values were the same, then it would be a horizontal line.